The following information relates to a firm for preparing a flexible budget: Column I II Details...

  1. The following information relates to a firm for preparing a flexible budget:

Column

I

II

Details

Amount ($)

Degree %

Repairs

1,300

60

Stores

1,000

100

Materials

16,179

100

Depreciation

5,000

nil

Wages

19,000

80

Advertising

2,349

75

Administration

1,450

25

Selling and distribution

3,000

70

Office Expenses

1,500

45

The amount for expenses is for production of 10,000 units. The amount changes with an increase or decrease of 20% above or below 10,000 units as indicated by the degree in column II above.

Required: show the budget statement
Answer #1
Units Degree of change
Expenses 8000 10000 12000 %
Repairs 1144 1300 1456 60
Stores 800 1000 1200 100
Materials 12943.2 16179 19414.8 100
Depreciation 5000 5000 5000 0
Wages 15960 19000 22040 80
Advertising 1996.65 2349 2701.35 75
Administration 1377.5 1450 1522.5 25
Selling and distribution 2580 3000 3420 70
Office Expenses 1365 1500 1635 45
Total Expenses 43166.35 50778 58389.65

Calculation Formulas in Excel :

Units Degree of change
Expenses =C2*(1-20%) 10000 =C2*(1+20%) %
Repairs =C3-(E3%*(C3*($C$2-$B$2)/$C$2)) 1300 =C3+(E3%*(C3*($D$2-$C$2)/$C$2)) 60
Stores =C4-(E4%*(C4*($C$2-$B$2)/$C$2)) 1000 =C4+(E4%*(C4*($D$2-$C$2)/$C$2)) 100
Materials =C5-(E5%*(C5*($C$2-$B$2)/$C$2)) 16179 =C5+(E5%*(C5*($D$2-$C$2)/$C$2)) 100
Depreciation =C6-(E6%*(C6*($C$2-$B$2)/$C$2)) 5000 =C6+(E6%*(C6*($D$2-$C$2)/$C$2)) 0
Wages =C7-(E7%*(C7*($C$2-$B$2)/$C$2)) 19000 =C7+(E7%*(C7*($D$2-$C$2)/$C$2)) 80
Advertising =C8-(E8%*(C8*($C$2-$B$2)/$C$2)) 2349 =C8+(E8%*(C8*($D$2-$C$2)/$C$2)) 75
Administration =C9-(E9%*(C9*($C$2-$B$2)/$C$2)) 1450 =C9+(E9%*(C9*($D$2-$C$2)/$C$2)) 25
Selling and distribution =C10-(E10%*(C10*($C$2-$B$2)/$C$2)) 3000 =C10+(E10%*(C10*($D$2-$C$2)/$C$2)) 70
Office Expenses =C11-(E11%*(C11*($C$2-$B$2)/$C$2)) 1500 =C11+(E11%*(C11*($D$2-$C$2)/$C$2)) 45
Total Expenses =SUM(B3:B12) =SUM(C3:C12) =SUM(D3:D12)